A venture capitalist, willing to invest $1,000,000, has three investments to choose from. The first investment, a software company, has a 10% chance of returning $5,000,000 profit, a 30% chance of returning $1,000,000 profit, and a 60% chance of losing the million dollars. The second company, a hardware company, has a 20% chance of returning $3,000,000 profit, a 40% chance of returning $1,000,000 profit, and a 40% chance of losing the million dollars. The third company, a biotech firm, has a 10% chance of returning $6,000,000 profit, a 70% of no profit or loss, and a 20% chance of losing the million dollars.
Construct a PDF for each investment.
Software company:
Hardware company:
Biotech firm:
Find the expected value for each investment.
software company:$
hardware company:$
biotech firm: $

Answer by ikleyn(53751)

.
A venture capitalist, willing to invest $1,000,000, has three investments to choose from.
The first investment, a software company, has a 10% chance of returning $5,000,000 profit,
a 30% chance of returning $1,000,000 profit, and a 60% chance of losing the million dollars.
The second company, a hardware company, has a 20% chance of returning $3,000,000 profit,
a 40% chance of returning $1,000,000 profit, and a 40% chance of losing the million dollars.
The third company, a biotech firm, has a 10% chance of returning $6,000,000 profit,
a 70% of no profit or loss, and a 20% chance of losing the million dollars.
(a) Construct a PDF for each investment.
Software company:
Hardware company:
Biotech firm:
(b) Find the expected value for each investment.
software company:$
hardware company:$
biotech firm: $
~~~~~~~~~~~~~~~~~~~~~~~~


        Part (a) is super-simple. Everybody, who is able to write or to print,
        can do it on his or her own without having any difficulties and without any help from outside.

        I will focus on part (b) here.


expected value, software company  E = 0.1*5000000 + 0.3*1000000 - 0.6*1000000 = 200,000 dollars.


expected value, hardware company  E = 0.2*3000000 + 0.4*1000000 - 0.4*1000000 = 600,000 dollars.


expected value, biotech  company  E = 0.1*6000000 + 0.7*0       - 0.2*2000000 = 200,000 dollars.

Solved.

As you see, this task is EXTREMELY simple and straightforward.
There is nothing on the way that would puzzle your mind.